To practice Problem-Solving Strategy 7.1 Rotational dynamics problems.

Suppose that you are holding a pencil balanced on its point. If you release the pencil and it begins to fall, what will be the angular acceleration when it has an angle of 10.0 degrees from the vertical?

D) Sort the forces as producing a torque of positive, negative, or zero magnitude about the rotational axis identified in Part A. Keep in mind that counterclockwise rotations are positive.

answer choices include normal force and weight. they have to be sorted into positive, negative, or zero.
Answer #1

Moment of Inertia of Long Thin Rod at end is I = m*(L^2) / 3
|r| |F| sin(?) = [m * (L^2) / 3] * ?
|L / 2| |m * g| sin(?) / [m * (L^2) / 3] = ?
3 g sin(?) / (2 * L) = ?

So....
3 (9.8) sin(10) / (2 * 0.15) = 17 rad/s^2 = ? *if accelerating counter clockwise*

If it is accelerating clockwise then ? = -17 rad/s^2